# Infrared Spectroradiometry of Lithium Benzoate from 21 to 235 THz

### v2

@@ Abstract
This paper presents an extensive survey of the thermal radiation properties of lithium benzoate. We heated the sample from 313 to 553 K, just below the melting point, while performing an infrared spectroradiometry with an FT-IR spectrometer from 21 to 235 THz (700-7800 cm-1). We have provided a detailed analysis of the infrared spectrum data and a comparison of the absorption spectrum of the same sample. It turned out that the recorded spectra are not only different from ordinary absorption spectra but also carry substantial information about the temperature dependence of the population of vibrationally excited states. We conclude by proposing a hypothesis on the thermal excitation mechanism of vibrational energy levels of molecules consistent with the distinct characteristics of the obtained infrared emission spectra.
@@ 1 Introduction
Despite a long history of infrared spectroradiometry, primarily in the steel industry (see e.g., [4, 5, 6, 7, 8, 9, 10]), most of the research focuses only on the mid-Infrared (mid-IR) region of the spectrum (400-4000 cm-1), or relatively high temperatures over 400 K. In this paper, we take a step forward and perform infrared spectroradiometry of an organic molecule, lithium benzoate, from 21 to 235 THz (700-7800 cm-1), encompassing both mid-IR and near-infrared (NIR) regions over 4000 cm-1. For a detailed investigation of the temperature dependence of the infrared (IR) emission of lithium benzoate, we systematically recorded its IR emission from 313 to 553 K in 40-degree increments. This focus helps to elucidate the temperature-dependent spectral features within the studied frequency range.
@@ 2.1 Experimental setup
Lithium benzoate (CAS No. 553-54-8) was purchased from BLD Pharm, Inc. The sample, a powder with a purity of 99%, was placed in a 10 mm diameter aluminum cup, forming a layer approximately 0.5 mm thick.
@@ 3.1 Overview of obtained spectra
Fig. 2-(a) shows the relative IR emission spectra of lithium benzoate. The overall spectral profiles in the mid-IR region exhibit little change with increasing temperature. In contrast, the NIR spectra become sharper at elevated temperatures. As an illustration, the spectra around 4600 cm-1 appear fuzzy at low temperatures, due to atmospheric CO2 absorption, but become clearer as the sample is heated. This trend suggests that enhanced molecular motion at higher temperatures increases the population of higher vibrational levels, leading to the emergence of more distinct NIR spectral features.
@@
Figure 2: (a) Temperature dependence of the relative IR emission spectra of lithium benzoate powder. (b) An enlargement of (a) into the region 24-36 THz (800-1190 cm-1). (c) The degree of shifts of all numbered peaks in (b) relative to those at the lowest temperature.
@@ 3.2 Details of mid-IR spectra
Figure 2-(c) shows the peak shifts relative to those in the lowest temperature, excluding the ninth peak. In general, all peaks broaden as the temperature increases. One of the causes of line broadening may be the shortening of the lifetime of the vibrationally excited states due to increased collisional broadening at higher temperatures. The first eight peaks are red shifted at higher temperatures. Remarkably, the amount of shift is only about 0.1% of the total wavenumber, which may be due to the robustness of our sample, lithium benzoate, with a high melting point. Moreover, the tenth peak is almost invariant, although it may be due to the strong adjacent band. The ninth peak gradually splits into two pieces, as shown in Fig. 2-(b), but both exhibit small redshifts at higher temperatures (see Fig. 3).
@@ 3.3 Comparison between mid-IR absorption and emission spectra
We also recorded the mid-IR absorption spectrum of lithium benzoate at 300 K to compare it with the emission spectrum at 313 K. The infrared absorption spectrum was recorded using an FT-IR spectrometer (FT/IR-4200, JASCO) equipped with a DLATGS (Deuterated L-Alanine Triglycine Sulfate) detector, with 90 accumulations at a resolution of 4.0 cm-1. We have prepared the sample as a KBr pellet containing 1.0 wt% of the analyte. Potassium bromide (KBr, 99%, CAS No. 7758-02-3) was purchased from Tokyo Chemical Industry (TCI).
@@
Figure 4: Emission and absorption spectrum of lithium benzoate. The black line corresponds to the emission spectrum at 313 K, while the red line represents the absorption spectrum at 300 K.
@@ 4 Conclusion
In this paper, we performed the infrared spectroradiometry of lithium benzoate from 21 to 235 THz (700-7800 cm-1) at various temperatures. The obtained spectra show strong temperature dependence in the NIR region. In particular, the NIR spectra become sharper as the sample is heated, implying an increase in the population of higher vibrationally excited states.
In differentiating the absorption and emission spectra of lithium benzoate, we demonstrated that the emission spectrum provides more information than the absorption spectrum. To explain this characteristic feature of the emission spectra, we proposed the cascade-like thermal excitation and spontaneous emission mechanism. However, the complicated structure of the thermal emission prevented us from identifying the origin of all peaks in the emission spectra. More extensive investigations are needed to confirm our hypothesis and to extract much more information from emission spectra.

### v3

# Infrared Spectroradiometry of Sodium Benzoate from 21 to 235 THz

@@ Abstract
This paper presents an extensive survey of the thermal radiation properties of sodium benzoate. We heated the sample from 313 to 553 K, just below the melting point, while performing an infrared spectroradiometry with an FT-IR spectrometer from 21 to 235 THz (700-7800 cm-1). We have provided a detailed analysis of the infrared spectrum data and a comparison of the absorption spectrum of the same sample. It turned out that the recorded spectra are not only different from ordinary absorption spectra but also carry substantial information about the temperature dependence of the population of vibrationally excited states. We conclude by proposing a hypothesis on the thermal excitation mechanism of vibrational energy levels of molecules consistent with the distinct characteristics of the obtained infrared emission spectra.
@@ 1 Introduction
Despite a long history of infrared spectroradiometry, primarily in the steel industry (see e.g., [4, 5, 6, 7, 8, 9, 10]), most of the research focuses only on the mid-Infrared (mid-IR) region of the spectrum (400-4000 cm-1), or relatively high temperatures over 400 K. In this paper, we take a step forward and perform infrared spectroradiometry of an organic molecule, sodium benzoate, from 21 to 235 THz (700-7800 cm-1), encompassing both mid-IR and near-infrared (NIR) regions over 4000 cm-1. For a detailed investigation of the temperature dependence of the infrared (IR) emission of sodium benzoate, we systematically recorded its IR emission from 313 to 553 K in 40-degree increments. This focus helps to elucidate the temperature-dependent spectral features within the studied frequency range.
@@ 2.1 Experimental setup
Sodium benzoate was purchased from BLD Pharm (Shanghai, China; catalog number BD136438-25g, lot number BGW609) under the product name lithium benzoate (stated purity 99%) and was used as received. During a related study, the elemental composition of this lot was examined by inductively coupled plasma atomic emission spectroscopy (ICP-AES) using an ICPE-9820 spectrometer (Shimadzu Corp.). The analysis showed Na = 15.9 wt% and Li below the detection limit, identifying the material unambiguously as sodium benzoate. All measurements reported in this work were performed on this material. The sample, a powder with a purity of 99%, was placed in a 10 mm diameter aluminum cup, forming a layer approximately 0.5 mm thick.
@@ 3.1 Overview of obtained spectra
Fig. 2-(a) shows the relative IR emission spectra of sodium benzoate. The overall spectral profiles in the mid-IR region exhibit little change with increasing temperature. In contrast, the NIR spectra become sharper at elevated temperatures. As an illustration, the spectra around 4600 cm-1 appear fuzzy at low temperatures, due to atmospheric CO2 absorption, but become clearer as the sample is heated. This trend suggests that enhanced molecular motion at higher temperatures increases the population of higher vibrational levels, leading to the emergence of more distinct NIR spectral features.
@@
Figure 2: (a) Temperature dependence of the relative IR emission spectra of sodium benzoate powder. (b) An enlargement of (a) into the region 24-36 THz (800-1190 cm-1). (c) The degree of shifts of all numbered peaks in (b) relative to those at the lowest temperature.
@@ 3.2 Details of mid-IR spectra
Figure 2-(c) shows the peak shifts relative to those in the lowest temperature, excluding the ninth peak. In general, all peaks broaden as the temperature increases. One of the causes of line broadening may be the shortening of the lifetime of the vibrationally excited states due to increased collisional broadening at higher temperatures. The first eight peaks are red shifted at higher temperatures. Remarkably, the amount of shift is only about 0.1% of the total wavenumber, which may be due to the robustness of our sample, sodium benzoate, with a high melting point. Moreover, the tenth peak is almost invariant, although it may be due to the strong adjacent band. The ninth peak gradually splits into two pieces, as shown in Fig. 2-(b), but both exhibit small redshifts at higher temperatures (see Fig. 3).
@@ 3.3 Comparison between mid-IR absorption and emission spectra
We also recorded the mid-IR absorption spectrum of sodium benzoate at 300 K to compare it with the emission spectrum at 313 K. The infrared absorption spectrum was recorded using an FT-IR spectrometer (FT/IR-4200, JASCO) equipped with a DLATGS (Deuterated L-Alanine Triglycine Sulfate) detector, with 90 accumulations at a resolution of 4.0 cm-1. We have prepared the sample as a KBr pellet containing 1.0 wt% of the analyte. Potassium bromide (KBr, 99%, CAS No. 7758-02-3) was purchased from Tokyo Chemical Industry (TCI).
@@
Figure 4: Emission and absorption spectrum of sodium benzoate. The black line corresponds to the emission spectrum at 313 K, while the red line represents the absorption spectrum at 300 K.
@@ 4 Conclusion
In this paper, we performed the infrared spectroradiometry of sodium benzoate from 21 to 235 THz (700-7800 cm-1) at various temperatures. The obtained spectra show strong temperature dependence in the NIR region. In particular, the NIR spectra become sharper as the sample is heated, implying an increase in the population of higher vibrationally excited states.
In differentiating the absorption and emission spectra of sodium benzoate, we demonstrated that the emission spectrum provides more information than the absorption spectrum. To explain this characteristic feature of the emission spectra, we proposed the cascade-like thermal excitation and spontaneous emission mechanism. However, the complicated structure of the thermal emission prevented us from identifying the origin of all peaks in the emission spectra. More extensive investigations are needed to confirm our hypothesis and to extract much more information from emission spectra.